# Against The Achilles’ Heel: A Survey on Red Teaming for Generative Models

### v1

@@ 8.4 Evaluation
Different from evaluating LLMs, evaluation for applications should consider the output from the LLMs in the system as well as the possible effects they applied after interacting with environments or executing tools. This makes evaluation much more complex and difficult. Current works (Naihin et al., 2023; Ruan et al., 2023; Yuan et al., 2024) conduct evaluations by analyzing the interaction trajectories of LLMs. R-judge (Yuan et al., 2024) focuses on evaluating the safety awareness of LLMs by taking LLMs to analyze interactions with already executed actions. However, this includes taking unsafe actions or executing unsafe tools. To resolve this problem, AgentMonitor (Naihin et al., 2023) utilizes an LLM to predict and prevent any harmful content before the action is executed, while ToolEmu (Ruan et al., 2023) uses an LLM as an emulator to emulate the execution result and an evaluator to analyze the harmfulness and helpfulness of that action.
9 Future Direction
9.1 Systematic Exploration
A vast amount of pretraining data from different domains, combined with the generalization capability of language models, creates a large susceptible area of safe risks. It is made evident by the exploitation of data patterns (Completion Compliance) and that of generalization (Generalisation Glide). Existing approaches are often isolated and ad-hoc, which limits their exploration range, while searches suffer from diversity problems. The problem can be mitigated by crowd-sourcing the effort of competition (TensorTrust (Toyer et al., 2023)) and developing systematic searchers.
9.2 Evaluation
There’s a need for comprehensive, realistic red-teaming datasets that accurately represent potential attacks and a benchmark that compares the effectiveness of attack methods and the safety of models with unified metrics.
Existing benchmarks offer covered varieties of risk categories, greatly facilitating research related to the safety of LLMs. However, the richness of benchmarks also brings new challenges. The lack of unified standard evaluation metrics makes it difficult to make a fair comparison between different benchmarks (Souly et al., 2024).
Data homogeneity between various benchmarks also hinders efficient evaluation using fewer test examples (Xu et al., 2023c). These issues call for future research to solve the following problems: (1) how to establish standard evaluation metrics and frameworks; and (2) how to sample a subset of these benchmarks to reach similar evaluation results as using the whole.
9.3 Defense
Given that many approaches can effectively compromise the safety training ingrained within AI systems and the anticipated evolution of AI capabilities ranging from synthesizing complex biological agents to controlling critical infrastructures, the risks associated with those vulnerabilities are profound. In many areas, people find effective jailbreak approaches but there is still a lack of defense methods. In addition to the traditional defense against interventions to pretrain and fine-tune data, the task of removing sensitive information directly from model weights through model editing methods is considered promising (Patil et al., 2023; Hasan et al., 2024b). Effective and generalizable defense methods will also be in demand in the future as the arms race between attack and defense of AI goes on. In this section, we discuss potential future directions of future defense methods.
Multilingual
Research has shown that it is challenging to ensure multilingual safety alignment (Deng et al., 2023d; Wang et al., 2023c; Yong et al., 2023; Shen et al., 2024b). However, as AI-based applications interact with users across diverse linguistic and cultural landscapes, it’s essential to develop and implement safety protocols that also robust to this variety. Addressing multilingual vulnerabilities enhances the global usability of AI and strengthens the defense against nuanced, language-specific attacks that could otherwise exploit linguistic gaps in safety measures.
Multi-modality
A significant limitation in current research is the disparity in effectively managing the combination of data in multiple modalities during attacks. Some strategies focus solely on one modality. For instance, in a multi-modal LM, altering images can significantly affect text generation if the interaction between text and vision is not adequately considered. This lack of comprehensive modality integration reduces effectiveness in executing and defending against multi-modal attacks.
Weights Manipulation
Recent advancements have unveiled vulnerabilities of LLMs when attacked by malicious actors who jailbreak AI systems by manipulating their model weights (Lermen et al., 2023; Zhan et al., 2023; Qi et al., 2023b; Chen et al., 2023b; Qi et al., 2023b). Highlighted by the seminal work of (Kinniment et al., 2024), the need for preemptive and robust defense mechanisms against subversive fine-tuning practices is evident. Interpreting how model weights in different layers lead to the success of such attacks (Subhash et al., 2023) can help us to design weight-targeted defense methods and navigate safety & helpfulness trade-offs, paving the way for a future in which AI’s advantages can be harnessed securely and responsibly.
9.4 LLM Applications
While some works (Naihin et al., 2023; Ruan et al., 2023; Yuan et al., 2024) have proposed benchmarks to evaluate LLM-based applications, some issues remain unsolved. First, LLM applications often have a huge action space, while most works often evaluate part of trajectories, making the evaluation insufficient. Second, the question of how to execute actions safely remains open. Using a simulator (Ruan et al., 2023) to generate action outputs may not be accurate enough and can not reflect the effects on the real-world environment. For defense, current works generally propose strategies corresponding to specific attacking methods (Liu et al., 2023i; Zhang et al., 2024e; Pedro et al., 2023), while applications can be attacked through any component of the system. An effective defense needs to be a systematic strategy that considers any part of applications without hurting the performance of the whole system.
@@ 10 Conclusion
This survey offers a comprehensive overview of the entire pipeline from attack methods to defense strategies, highlighting the vulnerabilities of LLM-based applications and the rise of multilingual and multi-modal threats. We introduce a novel taxonomy rooted in model capabilities to categorize attack strategies and frame the generation of attack prompts as search problems, revealing the design space for future attack methodologies. Lastly, we suggest several directions for future research and emphasize the importance of cross-disciplinary collaboration for the development of secure and ethical LLMs. Our aspiration is to steer the scholarly community towards the enhancement of GenAI’s reliability and trustworthiness, recognizing the pivotal role of robust, collaborative efforts in this endeavor.
Appendix
@@ Appendix A Full Paper List
Perez et al. (2022); Ganguli et al. (2022); Zhuo et al. (2023); Greshake et al. (2023); Xu et al. (2023b); Roy et al. (2023a); Yang et al. (2023d); Liu et al. (2023g); Xue et al. (2023); Casper et al. (2023); Qi et al. (2023a); Zou et al. (2023); Shayegani et al. (2023a); Zhang et al. (2023e); Bagdasaryan et al. (2023); Kandpal et al. (2023); Pedro et al. (2023); Shen et al. (2023); Yuan et al. (2023); Bhardwaj & Poria (2023b); Titus & Russell (2023); Schlarmann & Hein (2023); Yu et al. (2023a); Yao et al. (2023a); Lapid et al. (2023); Dong et al. (2023); Ma et al. (2023); Chin et al. (2023); Lermen et al. (2023); Yang et al. (2023b); Yong et al. (2023); Schulhoff et al. (2023); Xu et al. (2023e); Chao et al. (2023); Liu et al. (2023d); Zhang et al. (2023b); Huang et al. (2023b); Jiang et al. (2023d); Wei et al. (2023b); Gade et al. (2023); Zhu et al. (2023); Srivastava et al. (2023); Roy et al. (2023b); Bhardwaj & Poria (2023a); tse Huang et al. (2023); Tian et al. (2023b); Zhan et al. (2023); Li et al. (2023c); Xu et al. (2023d); Ding et al. (2023); Shah et al. (2023); Zhang et al. (2023c); Wang & Shu (2023); Wu et al. (2023c); Gong et al. (2023); Yang et al. (2023c); Mehrabi et al. (2023b); Qiang et al. (2023); Radharapu et al. (2023); Wang et al. (2023e); Fu et al. (2023a); Fishchuk & Braun (2023); Huang & Baldwin (2023); Cao et al. (2023b); Wen et al. (2023); Jiang et al. (2023a); Singh et al. (2023); Liu et al. (2023c); Rando & Tramèr (2023); Wang et al. (2023b); Mo et al. (2023); Du et al. (2023); Mehrotra et al. (2023); Fort (2023); Shahgir et al. (2023); He et al. (2023a); Wang et al. (2023d); Fu et al. (2023b); Pelrine & et al. (2023); Sheng et al. (2023); Salem et al. (2023); Zhang et al. (2023d); Jeong (2023); Oh et al. (2023); Wu et al. (2023a); Jiang et al. (2023c); Greenblatt et al. (2023); Takemoto (2024a); Zeng et al. (2024a); Zhang et al. (2024e); Wichers et al. (2024); Li et al. (2024c; a); Xiang et al. (2024); Hubinger et al. (2024); Zhao et al. (2024c); Salinas & Morstatter (2024); Ropers et al. (2024); Feffer et al. (2024); Guo et al. (2024); Chang et al. (2024); Lemkin (2024); Liu et al. (2024a; b); Zhang et al. (2024a); Wang et al. (2024a); Pang et al. (2024); Mangaokar et al. (2024); Sha & Zhang (2024); Shen et al. (2024a); Raina et al. (2024); Deng et al. (2024); Zhou et al. (2024d); Qiang et al. (2024); Lu et al. (2024); Zhang et al. (2024b); Zhao et al. (2024b); Xu et al. (2024a); Zou et al. (2024); Chu et al. (2024b); Qraitem et al. (2024); Lemkin (2024); Banerjee et al. (2024); Mo et al. (2024a); Wang et al. (2024e); Kim (2024); Geiping et al. (2024); Sadasivan et al. (2024); Li et al. (2024e); Gu et al. (2024); Sun & Miceli-Barone (2024); Pasquini et al. (2024); Li et al. (2024f); Fu et al. (2024); Li et al. (2024g); Xu & Wang (2024); Xiao et al. (2024)
@@
Sun et al. (2023); Aroyo et al. (2023); Xu et al. (2023a); Qiu et al. (2023); Wang et al. (2023g); Röttger et al. (2023); Ruan et al. (2023); Liu et al. (2023a); Zhang et al. (2023f); Wang et al. (2023c); Xu et al. (2023c); Mu et al. (2023); Tu et al. (2023); Liu et al. (2023e); Hung et al. (2023); Vidgen et al. (2023); Toyer et al. (2023); Huang et al. (2023a); Yi et al. (2023); Yuan et al. (2024); Li et al. (2024b); Souly et al. (2024); Mazeika et al. (2024)
Defense
Meade et al. (2023); Khalatbari et al. (2023); Tian et al. (2023a); Ji et al. (2023); Deng et al. (2023b); Alon & Kamfonas (2023); Phute et al. (2023); Bianchi et al. (2023); Kumar et al. (2023); Jain et al. (2023); Zhang et al. (2023h); Patil et al. (2023); Cao et al. (2023a); Qi et al. (2023b); Wang et al. (2023h); Dai et al. (2023); Rebedea et al. (2023); Robey et al. (2023); Chen et al. (2023a); Kim et al. (2023); Naihin et al. (2023); Shi et al. (2023); Pisano et al. (2023); Zhang et al. (2023g); Costa-jussà et al. (2023); Hu et al. (2023); Cui et al. (2023); He et al. (2023b); Inan et al. (2023); Varshney et al. (2023); Piet et al. (2023); Zhang et al. (2023d); Li et al. (2024d); Zhang et al. (2024d); Pi et al. (2024); Hasan et al. (2024a); Chern et al. (2024); Zhou et al. (2024a); Zheng et al. (2024); Kumar et al. (2024); Suo (2024); Chan et al. (2024); Wu et al. (2024); Sharma et al. (2024); Wang et al. (2024b); Ji et al. (2024); Rosati et al. (2024); Zhang et al. (2024f); Zhou et al. (2024c); Chen et al. (2024); Mo et al. (2024b); Yung et al. (2024); Xu et al. (2024b); Moraffah et al. (2024); Koh et al. (2024); Hua et al. (2024); Wang et al. (2024c); Kim et al. (2024); Zong et al. (2024); Wang et al. (2024d); Gou et al. (2024); Zeng et al. (2024b); Longpre et al. (2024); Hu et al. (2024); Han et al. (2024); Luo et al. (2024); Zhao et al. (2024a)
@@
Subhash (2023); Kang et al. (2023); Derner & Batistič (2023); Yu et al. (2023b); Zhang et al. (2023a); Inie et al. (2023); Wang et al. (2023f); Zhao et al. (2023); Antebi et al. (2024); Shen et al. (2024b); Wolf et al. (2024); Lee et al. (2024); Ye et al. (2024); Zhou et al. (2024e); Wei et al. (2024)
Survey
Deng et al. (2023c); Rao et al. (2023); Mozes et al. (2023); Shayegani et al. (2023b); Li et al. (2023a); Schwinn et al. (2023); Wang et al. (2023a; a); Esmradi et al. (2023); Rossi et al. (2024); Li et al. (2024h); Sun et al. (2024); Cui et al. (2024); Das et al. (2024); Xu et al. (2024c); Chu et al. (2024a); Chowdhury et al. (2024); Abdali et al. (2024); Pankajakshan et al. (2024)
Taxonomy
Kirk et al. (2023); Liu et al. (2023h); Park et al. (2023); Lambert et al. (2023); Derner et al. (2023)

### v2

@@ 8.4 LLM Applications Safety Evaluation
Different from evaluating LLMs, evaluation for applications should consider the output from the LLMs as well as the possible effects after interacting with the environment or executing tools. This makes evaluation difficult. Current work Naihin et al. (2023); Ruan et al. (2023); Yuan et al. (2024) conducts evaluation by analyzing the interaction trajectories of LLMs. R-judge Yuan et al. (2024) focuses on evaluating the safety awareness of LLMs by analyzing interactions with previously-executed actions. However, this includes taking unsafe actions or executing unsafe tools. To resolve this problem, AgentMonitor Naihin et al. (2023) uses an LLM to predict and prevent any harmful content before the action is executed, while ToolEmu Ruan et al. (2023) uses an LLM to emulate the execution result and an evaluator to analyze the harmfulness and helpfulness of that action.
9 Future Directions
As generative AI technology rapidly evolves, new risks emerge, particularly in cybersecurity, persuasive capabilities, privacy, and domain-specific applications. Addressing these challenges requires a diverse and realistic evaluation framework that moves beyond current benchmarks and detection methods. By incorporating adaptive, cross-cultural, and tool-integrated safety measures, evaluation methodologies can better account for emerging threats and usage contexts. In parallel, defensive strategies must evolve to address multilingual, multimodal, and model-weight manipulation threats, promoting resilience against sophisticated attacks. In this section, we outline key directions for generative AI safety research.
9.1 Expanding the Safety Landscape
Extensive research has investigated the safety concerns associated with language models, with early efforts primarily focused on biases Kiritchenko & Mohammad (2018); Han et al. (2021). As generative AI has progressed, recent work has shifted toward addressing the generation of harmful content, such as instructions for dangerous or unethical behavior Wang et al. (2023g); Mazeika et al. (2024). However, new safety concerns continue to emerge, which we identify as promising directions for future investigation.
  • Cybersecurity: LLMs have demonstrated impressive capabilities in generating functional code from natural language requests Li et al. (2023d); Liu et al. (2023j). However, relatively little work has focused on cybersecurity, where risks are particularly significant. Key examples include Pearce et al. (2022), which uses handcrafted prompts based on the top 25 MITRE Common Weakness Enumeration (CWE) entries in Python, C, and Verilog; Siddiq & Santos (2022), covering 75 CWEs in Python; and Bhatt et al. (2023), which extends evaluation to four additional programming languages. These studies share common limitations, such as single-turn queries, English-only test cases, and imperfect detection of insecure coding patterns. Future work could focus on expanding the work to address these limitations.
  • LLM Deception, Flattery, and Persuasion: Recent work has indicated that LLMs may exhibit deceptive Yao et al. (2024); Anonymous (2024), flattering (sycophantic) Ranaldi & Pucci (2023), or persuasive behaviors Salvi et al. (2024). While these behaviors are not always harmful, they are associated with various risks. Despite these potential concerns, few studies have explored mitigation strategies or evaluated these behaviors as safety issues Pacchiardi et al. (2024); Chen et al. (2024b). Given the growing use of LLMs by general audiences, these issues warrant increased attention and further research.
  • Privacy Violations: Although numerous studies have addressed privacy issues in generative AI, research on privacy leakage remains limited, often relying on synthetic data or simplistic attacks in controlled settings Chen et al. (2023b); Gupta et al. (2023). More realistic and empirical evaluations of privacy risks are necessary Li et al. (2023a).
  • Domain-Specific Risks: The vast diversity of LLM pretraining data, coupled with their generalization abilities, introduces vulnerabilities in various domains (recall Section 4.1.3). While research has primarily focused on high-stakes areas such as finance, medicine, and law Chen et al. (2024d), further investigation into low-resource or other high-risk domains (e.g., chemical, nuclear) is essential.
  • Safety in Real-time Interaction or LLM Agents: LLMs are increasingly central to agent systems. Current research on LLM-based application safety Deng et al. (2024); Yang et al. (2024b); Gu et al. (2024); Ruan et al. (2023) mainly examines limited scenarios with specific agents in simulated environments. Future research should consider more dynamic, real-time environments where LLMs interact with additional tools, hold higher privileges, or function with enhanced autonomy.
9.2 Unified and Realistic Evaluation
In Section 6, we reviewed various evaluation metrics and benchmarks. However, many issues and limitations remain. We propose several directions for advancing LLM safety evaluation.
  • Diverse and Dynamic Benchmarks: Existing benchmarks often exhibit high data homogeneity. For instance, benchmarks like DoNotAnswer Wang et al. (2023g), AdvBench Zou et al. (2023b), and HarmfulQ Shaikh et al. (2023) are highly similar to one another, although DoNotAnswer offers broader risk coverage. This limits the efficiency of model evaluation with fewer test examples Xu et al. (2023c). Future research on benchmarks should focus on: (1) covering a wider range of safety areas as discussed in Section 9.1; (2) creating benchmarks that evolve dynamically over time to incorporate and assess emerging risks; (3) expanding safety evaluations to cover underrepresented languages and cultures; and (4) developing benchmarks for tool and API integration safety.
  • Standardized Metrics and Detection Methods: Many studies employ metrics like ASR, but the definition and detection of “success” in ASR varies. Common detection methods include: (1) string matching of affirmative terms like “Sure,”; (2) classification by a fine-tuned small model; and (3) using LLMs to evaluate harmful content through prompts. This variation leads to inconsistent comparisons across benchmarks, as unified evaluation standards are lacking Souly et al. (2024), and some methods can produce false positives. Establishing standardized metrics and unified implementations is critical for comparability and reproducibility.
  • Balancing Safety and Helpfulness: Some studies have indicated that increasing safety can decrease general model performance Shi et al. (2023b); Zhang et al. (2024c); Cui et al. (2024a). Research on mitigating this trade-off or finding an optimal balance between safety and helpfulness is a promising area. We also advocate for future safety research to incorporate not only safety-specific evaluation but also general assessments that offer a more comprehensive understanding of model behavior.
  • LLM-based Evaluators: LLMs have shown utility as evaluators in many studies Zheng et al. (2023b); Wang et al. (2023g); Li et al. (2024a). However, commercial models are often overly cautious, rejecting most requests with even moderately risky content, which presents challenges for using LLMs as safety judges. Few studies have explored fine-tuning an LLM to act specifically as a safety evaluator. Future research could focus on optimizing LLM-based evaluation methods to enhance effectiveness in safety assessment.
9.3 Advanced Defense Mechanisms
As generative AI evolves rapidly, future defense research must address emerging vulnerabilities, including linguistic diversity, multimodality, adversarial attacks, and model-weight manipulation. Key future directions are outlined below.
  • Multilingual Defense: As introduced in Section 4.1.3, multilingual attacks have been recognized as a simple and effective method Deng et al. (2023d); Wang et al. (2023c); Yong et al. (2023); Shen et al. (2024b). Effective defense in multilingual contexts is challenging, as each language presents unique vulnerabilities. Developing multilingual safety alignment techniques that address language-specific structures and cultural contexts is a promising approach to mitigating language-specific attacks.
  • Multimodality Defense: The safety research landscape often lags behind model development. With recent advancements, multimodal LLMs are now capable of processing diverse inputs, and more teams are involved in developing these models. However, safety research in multimodal LLMs currently tends to focus on limited scenarios, such as controlling for explicit images or toxic text generation, while overlooking cross-modal attacks He et al. (2023a); Shayegani et al. (2023a) and defense in audio Xu et al. (2024b) and video modalities Wang et al. (2024g). Future research should expand to include these areas, with improved defense methods and ethical considerations for responsible multimodal LLM usage.
  • Defense Against Weight Manipulation: Recent findings reveal vulnerabilities introduced through malicious fine-tuning of model weights Lermen et al. (2023); Zhan et al. (2023); Qi et al. (2023); Chen et al. (2023b); Qi et al. (2023). Studies like Kinniment et al. (2023) highlight the need for robust, preemptive defenses against such practices. Understanding how weights in different layers contribute to the success of these attacks Subhash et al. (2023) could aid in developing resilient, weight-targeted defense strategies while balancing model utility. Additional research that provides analysis and explanatory studies in this area would further benefit the field and represents a valuable direction for future work.
  • Defense Strategies through Model Manipulation: Many models have been pretrained on sensitive data that cannot easily be removed from model weights, posing ongoing risks. Removing sensitive information directly from weights has become a critical research area. Techniques such as selective pruning, structured editing, and targeted weight adjustments have shown promise in initial work Patil et al. (2023); Hasan et al. (2024b). Continued research into these and other novel approaches for precise content retention control is essential for future work.
@@ 10 Conclusion
This survey offers a comprehensive overview of the entire pipeline from attack methods to defense strategies, highlighting the vulnerabilities of LLM-based applications and the rise of multilingual and multimodal threats. We introduce a novel taxonomy rooted in model capabilities to categorize attack strategies and frame the generation of attack prompts as search problems, revealing the design space for future attack methodologies. Lastly, we suggest several directions for future research and emphasize the importance of cross-disciplinary collaboration for the development of secure and ethical LLMs. Our aspiration is to steer the scholarly community toward the enhancement of GenAI’s reliability and trustworthiness, recognizing the pivotal role of robust, collaborative efforts in this endeavor.
@@ Appendix A Full Paper List
Survey
Deng et al. (2023c); Rao et al. (2023); Mozes et al. (2023); Shayegani et al. (2023b); Li et al. (2023a); Schwinn et al. (2023); Wang et al. (2023a; a); Esmradi et al. (2023); Rossi et al. (2024); Li et al. (2024j); Sun et al. (2024); Cui et al. (2024b); Das et al. (2024); Xu et al. (2024f); Chu et al. (2024a); Chowdhury et al. (2024); Abdali et al. (2024); Pankajakshan et al. (2024)
Taxonomy
Kirk et al. (2023a); Liu et al. (2023h); Park et al. (2023); Lambert et al. (2023); Derner et al. (2023)
@@
Perez et al. (2022); Ganguli et al. (2022); Zhuo et al. (2023); Abdelnabi et al. (2023); Xu et al. (2023b); Roy et al. (2023a); Yang et al. (2023d); Liu et al. (2023g); Xue et al. (2023); Casper et al. (2023); Qi et al. (2024); Zou et al. (2023b); Shayegani et al. (2023a); Zhang et al. (2023e); Bagdasaryan et al. (2023); Kandpal et al. (2023); Pedro et al. (2023); Shen et al. (2023); Yuan et al. (2023b); Bhardwaj & Poria (2023b); Titus & Russell (2023); Schlarmann & Hein (2023); Yu et al. (2023a); Yao et al. (2023a); Lapid et al. (2023); Dong et al. (2023); Ma et al. (2023); Chin et al. (2023); Lermen et al. (2023); Yang et al. (2023b); Yong et al. (2023); Schulhoff et al. (2023); Xu et al. (2023e); Chao et al. (2023); Liu et al. (2023d); Zhang et al. (2023b); Huang et al. (2023c); Jiang et al. (2023d); Wei et al. (2023b); Gade et al. (2023); Zhu et al. (2023); Srivastava et al. (2023); Roy et al. (2023b); Bhardwaj & Poria (2023a); Huang et al. (2023a); Tian et al. (2023b); Zhan et al. (2023); Li et al. (2023e); Xu et al. (2023d); Ding et al. (2023); Shah et al. (2023); Zhang et al. (2023c); Wang & Shu (2023); Wu et al. (2023d); Gong et al. (2023); Yang et al. (2023c); Mehrabi et al. (2023b); Qiang et al. (2023); Radharapu et al. (2023); Wang et al. (2023e); Fu et al. (2023a); Fishchuk & Braun (2023); Huang & Baldwin (2023); Cao et al. (2023b); Wen et al. (2023); Jiang et al. (2023a); Singh et al. (2023); Liu et al. (2023c); Rando & Tramèr (2023); Wang et al. (2023b); Mo et al. (2023); Du et al. (2023b); Mehrotra et al. (2023); Fort (2023); Shahgir et al. (2023); He et al. (2023a); Wang et al. (2023d); Fu et al. (2023b); Pelrine & et al. (2023); Sheng et al. (2023); Salem et al. (2023); Zhang et al. (2023d); Jeong (2023); Oh et al. (2023); Wu et al. (2023b); Jiang et al. (2023c); Greenblatt et al. (2023); Takemoto (2024a); Zeng et al. (2024a); Zhang et al. (2024e); Wichers et al. (2024); Li et al. (2024d; b); Xiang et al. (2024); Hubinger et al. (2024); Zhao et al. (2024c); Salinas & Morstatter (2024); Ropers et al. (2024); Feffer et al. (2024); Guo et al. (2024); Chang et al. (2024); Lemkin (2024); Liu et al. (2024a; b); Zhang et al. (2024a); Wang et al. (2024a); Pang et al. (2024); Mangaokar et al. (2024); Sha & Zhang (2024); Shen et al. (2024a); Raina et al. (2024); Deng et al. (2024); Zhou et al. (2024d); Qiang et al. (2024); Lu et al. (2024a); Zhang et al. (2024b); Zhao et al. (2024b); Xu et al. (2024c); Zou et al. (2024b); Chu et al. (2024b); Qraitem et al. (2024); Lemkin (2024); Banerjee et al. (2024); Mo et al. (2024a); Wang et al. (2024h); Kim (2024); Geiping et al. (2024); Sadasivan et al. (2024); Li et al. (2024g); Gu et al. (2024); Sun & Barone (2024); Pasquini et al. (2024); Li et al. (2024h); Fu et al. (2024); Li et al. (2024i); Xu & Wang (2024); Xiao et al. (2024)
Defense
Meade et al. (2023); Khalatbari et al. (2023); Tian et al. (2023a); Ji et al. (2023); Deng et al. (2023b); Alon & Kamfonas (2023); Phute et al. (2023); Bianchi et al. (2023); Kumar et al. (2023); Jain et al. (2023); Zhang et al. (2023h); Patil et al. (2023); Cao et al. (2023a); Qi et al. (2023); Wang et al. (2023h); Dai et al. (2023); Rebedea et al. (2023); Robey et al. (2023); Chen et al. (2023a); Kim et al. (2023); Naihin et al. (2023); Shi et al. (2023b); Pisano et al. (2023); Zhang et al. (2023g); Costa-jussà et al. (2023); Hu et al. (2023); Cui et al. (2023); He et al. (2023b); Inan et al. (2023); Varshney et al. (2024); Piet et al. (2023); Zhang et al. (2023d); Li et al. (2024f); Zhang et al. (2024d); Pi et al. (2024); Hasan et al. (2024a); Chern et al. (2024); Zhou et al. (2024a); Zheng et al. (2024); Kumar et al. (2024); Suo (2024); Chan et al. (2024); Wu et al. (2024); Sharma et al. (2024); Wang et al. (2024b); Ji et al. (2024); Rosati et al. (2024); Zhang et al. (2024f); Zhou et al. (2024c); Chen et al. (2024a); Mo et al. (2024b); Yung et al. (2024); Xu et al. (2024d); Moraffah et al. (2024); Koh et al. (2024); Hua et al. (2024); Wang et al. (2024d); Kim et al. (2024); Zong et al. (2024); Wang et al. (2024e); Gou et al. (2024); Zeng et al. (2024b); Longpre et al. (2024); Hu et al. (2024); Han et al. (2024); Luo et al. (2024); Zhao et al. (2024a)
@@
Sun et al. (2023); Aroyo et al. (2023); Xu et al. (2023a); Qiu et al. (2023); Wang et al. (2023g); Röttger et al. (2023); Ruan et al. (2023); Liu et al. (2023a); Zhang et al. (2023f); Wang et al. (2023c); Xu et al. (2023c); Mu et al. (2023); Tu et al. (2023); Liu et al. (2023e); Hung et al. (2023); Vidgen et al. (2023); Toyer et al. (2023); Huang et al. (2023b); Yi et al. (2023); Yuan et al. (2024); Li et al. (2024c); Souly et al. (2024); Mazeika et al. (2024)
@@
Subhash (2023); Kang et al. (2023); Derner & Batistic (2023); Yu et al. (2023b); Zhang et al. (2023a); Inie et al. (2023); Wang et al. (2023f); Zhao et al. (2023); Antebi et al. (2024); Shen et al. (2024b); Wolf et al. (2024); Lee et al. (2024); Ye et al. (2024); Zhou et al. (2024e); Wei et al. (2024)